# Learning for Dynamic Combinatorial Optimization without Training Data

### v1

@@ Abstract
We introduce DyCO-GNN, a novel unsupervised learning framework for Dynamic Combinatorial Optimization that requires no training data beyond the problem instance itself. DyCO-GNN leverages structural similarities across time-evolving graph snapshots to accelerate optimization while maintaining solution quality. We evaluate DyCO-GNN on dynamic maximum cut, maximum independent set, and the traveling salesman problem across diverse datasets of varying sizes, demonstrating its superior performance under tight and moderate time budgets. DyCO-GNN consistently outperforms the baseline methods, achieving high-quality solutions up to 3–60x faster, highlighting its practical effectiveness in rapidly evolving resource-constrained settings.
@@ 1 Introduction
Combinatorial optimization (CO) lies at the heart of many critical scientific and industrial problems [37]. Since most CO problems are NP-hard, solving large-scale instances is computationally prohibitive. Traditionally, solving such problems has relied on exact solvers, heuristics, and metaheuristics, whose design requires problem-specific insights. Recent advances in machine learning, particularly graph neural networks (GNNs), have opened new avenues for learning heuristics in a data-driven manner [23, 24, 10, 18, 25, 6, 26]. GNNs have emerged as a powerful framework for learning over relational and structured data [7, 28, 15, 49], with an inductive bias particularly suited for representing the underlying graph structures inherent in many CO problems.
Most existing approaches applying machine learning to CO are learning-heavy, requiring training on a large set of problem instances to learn heuristics that generalize to test instances. Training can take hours or even days on multiple GPUs [51, 33]. Schuetz et al. [42] started another line of research that aims to develop an unsupervised learning (UL) method that learns instance-specific heuristics. Their method, PI-GNN, directly applies a GNN to the problem instance of interest and optimizes a CO objective. Under this design, no explicit “training” is involved; the runtime is the time it takes for the model to converge on each problem instance. Prior work has shown that problems such as maximum cut (MaxCut) and maximum independent set (MIS) on graphs with thousands of nodes can be solved by PI-GNN-based methods in minutes [17, 19]. We provide further runtime analysis in Section 5.
While significant progress has been made in learning for static CO, many real-world problems are inherently dynamic, involving inputs or constraints that evolve over time [53, 55]. In such settings, decisions must be updated continually, and practical algorithms must be efficient. A simple approach is to treat each snapshot of the problem instance as a static problem and solve it from scratch. However, significant overlap often exists in the node and edge sets across snapshots. Leveraging information from previous snapshots can improve both runtime and solution quality. For example, in MaxCut, the previous solution might serve as a good starting point after edge additions.
In this paper, we focus on learning for dynamic combinatorial optimization (DCO) and aim to advance the PI-GNN line of research. We present DyCO-GNN for DCO. To the best of our knowledge, our work is the first to apply machine learning to DCO problems. Our main contributions are as follows.
@@ 2 Related work
Machine learning for static CO has predominantly relied on supervised learning approaches [23, 24, 50, 10, 44, 18, 32, 33]. These methods train models to predict high-quality solutions given a large dataset of problem instances paired with optimal or near-optimal labels. However, generating such labels is computationally expensive, especially for large-scale instances, making these methods less practical for many real-world applications.
To address this limitation, UL and reinforcement learning (RL) approaches have emerged as promising alternatives [6, 26, 29, 25, 39, 47, 48, 51, 41]. These methods aim to learn optimization strategies directly from instance structures or reward signals without requiring labeled data. While they alleviate the need for ground-truth solutions, most UL and RL-based techniques still involve extensive offline training over large datasets to develop heuristics that generalize. Training such models can be time-consuming, often requiring hours or days on high-performance hardware [51, 33].
A different paradigm was introduced by Schuetz et al. [42], who proposed PI-GNN—an unsupervised framework that learns instance-specific heuristics by directly optimizing the CO objective on a single instance. This avoids any offline training and allows the method to be tailored to each problem instance during inference. PI-GNN combines a learnable embedding layer with a GNN and has been shown to achieve competitive performance across tasks such as MaxCut and MIS. Subsequent works have improved its solution quality and extended its design to incorporate higher-order relational reasoning [17, 19], achieving strong results even on large-scale graph instances.
@@ 3.1 Static CO: PI-GNN and quadratic binary unconstrained optimization (QUBO)
Consider a graph $G_{p}=(V_{p},E_{p},w_{p})$ with node set $V_{p}=\{1,2,\ldots,n_{p}\}$, edge set $E_{p}$, and edge weights $w_{p}$ representing or inherent in a CO problem instance $p$. PI-GNN is a general UL framework for static CO problems based on QUBO [34, 12, 8]. A QUBO problem is defined by:
$\min_{x}~{}~{}\ell(x;Q_{p})=x^{T}Q_{p}x=\sum_{i=1}^{N}\sum_{j=1}^{N}Q_{ij}x_{i}x_{j},$ (1)
@@
MaxCut. Without loss of generality, we assume the edges have unit weights. The goal of MaxCut is to find a partition of the nodes into two disjoint subsets that maximizes the number of edges between them. We can model MaxCut solutions using $x\in\{0,1\}^{n}$ where $x_{i}=0$ if node $i$ is in one set, and $x_{i}=1$ if node $i$ is in the other set. The objective can be formulated as $\sum_{(i,j)\in E}(2x_{i}x_{j}-x_{i}-x_{j})$, which is an instance of QUBO [12].
MIS. MIS is the largest subset of non-adjacent nodes. Similar to MaxCut, we let $x_{i}=1$ if node $i$ is in the independent set and $x_{i}=0$ otherwise. The objective can be formulated as $-\sum_{i\in V}x_{i}+M\sum_{(i,j)\in E}x_{i}x_{j}$, where $M\in\mathbb{R}_{>0}$ is a penalty term enforcing the independent set constraint [8].
TSP. We consider symmetric TSP instances modeled as undirected complete graphs. TSP aims to find the route that visits each node exactly once and returns to the origin node with the lowest total distance traveled. In contrast to MaxCut and MIS, we use a binary matrix $X\in\{0,1\}^{n\times n}$ to represent a route where $X_{ij}=1$ denotes visiting node $i$ at step $j$. Following [34, 20], we adopt the following objective: $\sum_{(i,j)\in E}w_{ij}\sum_{v=1}^{n}X_{iv}X_{j(v+1)}+M\sum_{i=1}^{n}(1-\sum_{j=1}^{n}X_{ij})^{2}+M\sum_{j=1}^{n}(1-\sum_{i=1}^{n}X_{ij})^{2}$, where $w_{ij}$ is the distance between node $i$ and node $j$, $X_{j(n+1)}=X_{j1}$, accounting for the requirement of returning to the origin node, and $M\in\mathbb{R}_{>0}$ is a penalty term to ensure $X$ represents a valid route. During the QUBO objective computation, we simply flatten $X$ to an $n^{2}$-dimensional vector.
@@ 3.2 DCO on graphs
We focus on DCO problems on discrete-time dynamic graphs (DTDGs). A DTDG representing a DCO instance is a sequence $[G_{p}^{(1)},G_{p}^{(2)},\ldots,G_{p}^{(T)}]$ of graph snapshots where each $G_{p}^{(t)}=(V_{p}^{(t)},E_{p}^{(t)},w_{p}^{(t)})$ has a node set $V_{p}^{(t)}=\{1,2,\ldots,n_{p}^{(t)}\}$, an edge set $E_{p}^{(t)}$, and edge weights $w_{p}^{(t)}$. For dynamic MaxCut and MIS, we consider the scenario where the node set $V_{p}^{(t)}$ and edge set $E_{p}^{(t)}$ change over time, and $w_{p}^{(t)}=\mathbf{1}$ for all $t\in\{1,2,\ldots,T\}$. For dynamic TSP, we consider the case where one of the nodes moves along a trajectory. Effectively, $V_{p}^{(1)}=V_{p}^{(2)}=\ldots=V_{p}^{(T)}$ and $E_{p}^{(1)}=E_{p}^{(2)}=\ldots=E_{p}^{(T)}$, but $w_{p}^{(t)}$ varies for each $t$.
Let $\Omega_{p}^{(t)}$ denote the set of discrete feasible solutions for snapshot $G_{p}^{(t)}$. For MaxCut and MIS, $\Omega_{p}^{(t)}$ is a subset of $\{0,1\}^{n_{p}^{(t)}}$. For TSP, $\Omega_{p}^{(t)}$ is the set of all possible routes that visit each node exactly once and return to the origin node. The objective is to find the optimal solution for each snapshot of a given DCO problem instance: $x^{(t)*}=\operatorname*{argmin}_{x\in\Omega_{p}^{(t)}}\ell(x;Q_{p}^{(t)})$, where $\ell(\cdot;Q_{p}^{(t)})$ denotes the cost function as defined earlier for each snapshot. For evaluation, we compute the mean approximation ratio (ApR) across all snapshots: Mean $\text{Mean ApR}=\frac{1}{T}\sum_{t=1}^{T}\ell(x^{(t)};Q_{p}^{(t)})/\ell(x^{(t)*};Q_{p}^{(t)})$.
@@ 4.1 Warm-starting PI-GNN
Given a sequence of graph snapshots $[G_{p}^{(1)},G_{p}^{(2)},\ldots,G_{p}^{(T)}]$, PI-GNN independently initiates an optimization process for each snapshot. One potential issue is that the time interval between two consecutive snapshots may be shorter than the convergence time of PI-GNN. In such cases, faster convergence is required. Considering the structural similarity across snapshots, a straightforward baseline method is to warm start the optimization using the parameters from the previous snapshot. Effectively, the “solution” to the previous snapshot serves as the initialization for the current one, allowing the model to “fine-tune” the solution instead of optimizing everything from scratch. We detail the procedure of warm-starting PI-GNN under different time budgets in Algorithm 1.
@@
2:Problem instance: $[G_{p}^{(1)},G_{p}^{(2)},\ldots,G_{p}^{(T)}]$
@@
6:Randomly initialize $\theta^{(1)}$
7:Construct $[Q_{p}^{(1)},Q_{p}^{(2)},\ldots,Q_{p}^{(T)}]$
8:for $[G_{p}^{(1)},G_{p}^{(2)},\ldots,G_{p}^{(T)}]$
@@
10:     Predict $x^{(1)}$ via $\mathcal{A}(G_{p}^{(1)};\theta^{(1)})$
11:     Update $\theta^{(1)}$ using $\nabla\ell(x^{(1)};Q_{p}^{(1)})$
@@
14:     $\theta^{(t)}\leftarrow\theta^{(t-1)}$
@@
16:         Predict $x^{(t)}$ via $\mathcal{A}(G_{p}^{(t)};\theta^{(t)})$
17:         Update $\theta^{(t)}$ using $\nabla\ell(x^{(t)};Q_{p}^{(t)})$
@@
2:Problem instance: $[G_{p}^{(1)},G_{p}^{(2)},\ldots,G_{p}^{(T)}]$
@@
6:Randomly initialize $\theta^{(1)}$
7:Construct $[Q_{p}^{(1)},Q_{p}^{(2)},\ldots,Q_{p}^{(T)}]$
8:for $[G_{p}^{(1)},G_{p}^{(2)},\ldots,G_{p}^{(T)}]$
@@
10:     Predict $x^{(1)}$ via $\mathcal{A}(G_{p}^{(1)};\theta^{(1)})$
11:     Update $\theta^{(1)}$ using $\nabla\ell(x^{(1)};Q_{p}^{(1)})$
@@
14:     $\theta^{(t)}\leftarrow\lambda_{\text{shrink}}\theta^{(t-1)}+\lambda_{\text{perturb}}\epsilon^{(t)}$
@@
16:         Predict $x^{(t)}$ via $\mathcal{A}(G_{p}^{(t)};\theta^{(t)})$
17:         Update $\theta^{(t)}$ using $\nabla\ell(x^{(t)};Q_{p}^{(t)})$
@@
Figure 1: Performance of static PI-GNN and warm-started PI-GNN on dynamic MaxCut and MIS instances. Detailed setup is explained in Section 5.
These phenomena are closely tied to the role of initialization in deep learning [45, 11, 16] and can be attributed to the highly nonconvex nature of the optimization objective with respect to the model parameters, which gives rise to numerous local optima. When the first snapshot is optimized to convergence, the model undergoes extensive optimization, resulting in increased confidence in its predictions. Consequently, when the model is warm-started, the gradients become small due to this overconfidence, making it more difficult for the model to escape local optima. This hampers its ability to discover higher-quality solutions despite faster initial progress.

### v2

@@ Abstract
We introduce DyCO-GNN, an unsupervised learning framework for Dynamic Combinatorial Optimization that requires no training data beyond the problem instance itself. DyCO-GNN leverages structural similarities across time-evolving graph snapshots to accelerate optimization while maintaining solution quality. We evaluate DyCO-GNN on dynamic maximum cut, maximum independent set, and the traveling salesman problem across diverse datasets of varying sizes, demonstrating its superior performance under tight and moderate time budgets. DyCO-GNN consistently outperforms the baseline methods, achieving high-quality solutions up to 3–60x faster, highlighting its practical effectiveness in rapidly evolving resource-constrained settings.
@@ 1 Introduction
Combinatorial optimization (CO) lies at the heart of many critical scientific and industrial problems (Papadimitriou and Steiglitz, 1982). Since most CO problems are NP-hard, solving large-scale instances is computationally prohibitive. Traditionally, solving such problems has relied on exact solvers, heuristics, and metaheuristics, whose design requires problem-specific insights. Recent advances in machine learning, particularly graph neural networks (GNNs), have opened new avenues for learning heuristics in a data-driven manner (Joshi et al., 2019; 2022; Gasse et al., 2019; Hudson et al., 2022; Karalias and Loukas, 2020; Bello et al., 2016; Khalil et al., 2017). GNNs have emerged as a powerful framework for learning over relational and structured data (Defferrard et al., 2016; Kipf and Welling, 2017; Hamilton et al., 2017; veličković2018graph), with an inductive bias particularly suited for representing the underlying graph structures inherent in many CO problems.
Most existing approaches applying machine learning to CO are learning-heavy, requiring training on a large set of problem instances to learn heuristics that generalize to test instances. Training can take hours or even days on multiple GPUs (Wang and Li, 2023; Li et al., 2024). Schuetz et al. (2022) started another line of research that aims to develop an unsupervised learning (UL) method that learns instance-specific heuristics. Their method, PI-GNN, directly applies a GNN to the problem instance of interest and optimizes a CO objective. Under this design, no explicit “training” is involved; the runtime is the time it takes for the model to converge on each problem instance. Prior work has shown that problems such as maximum cut (MaxCut) and maximum independent set (MIS) on graphs with thousands of nodes can be solved by PI-GNN-based methods in minutes (Heydaribeni et al., 2024; Ichikawa, 2024).
While significant progress has been made in learning for static CO, many real-world problems are inherently dynamic, involving inputs or constraints that evolve over time (Yang et al., 2012; Zhang et al., 2021). In such settings, decisions must be updated continually, and practical algorithms must be efficient. A simple approach is to treat each snapshot of the problem instance as a static problem and solve it from scratch. However, significant overlap often exists in the node and edge sets across snapshots. Leveraging information from previous snapshots can improve both runtime and solution quality. For example, in MaxCut, the previous solution might serve as a good starting point after edge additions. Our proposed approach leverages such overlaps.
Specifically, we focus on learning for dynamic combinatorial optimization (DCO) and aim to advance the PI-GNN line of research (i.e., instance-specific optimization). To the best of our knowledge, our work is the first general UL-based framework for DCO problems. Our main contributions are as follows.
@@ 2 Related work
Machine learning for static CO has predominantly relied on supervised learning approaches (Joshi et al., 2019; 2022; Vinyals et al., 2015; Gasse et al., 2019; Sun and Yang, 2023; Hudson et al., 2022; Li et al., 2023; 2024). These methods train models to predict high-quality solutions given a large dataset of problem instances paired with optimal or near-optimal labels. However, generating such labels is computationally expensive, especially for large-scale instances, making these methods less practical for many real-world applications.
To address this limitation, UL and reinforcement learning (RL) approaches have emerged as promising alternatives (Bello et al., 2016; Khalil et al., 2017; Kool et al., 2019; Karalias and Loukas, 2020; Qiu et al., 2022; Toenshoff et al., 2021; Tönshoff et al., 2023; Wang and Li, 2023; Sanokowski et al., 2023). These methods aim to learn optimization strategies directly from instance structures or reward signals without requiring labeled data. While they alleviate the need for ground-truth solutions, most UL and RL-based techniques still involve extensive offline training over large datasets to develop heuristics that generalize. Training such models can be time-consuming, often requiring hours or days on high-performance hardware (Wang and Li, 2023; Li et al., 2024).
A different paradigm was introduced by Schuetz et al. (2022), who proposed PI-GNN—an unsupervised framework that learns instance-specific heuristics by directly optimizing the CO objective on a single instance. This avoids any offline training and allows the method to be tailored to each problem instance during inference. PI-GNN combines a learnable embedding layer with a GNN and has been shown to achieve competitive performance across tasks such as MaxCut and MIS. Subsequent works have improved its solution quality and extended its design to incorporate higher-order relational reasoning (Heydaribeni et al., 2024; Ichikawa, 2024), achieving strong results even on large-scale graph instances.
@@ 3.1 Static CO: PI-GNN and quadratic binary unconstrained optimization (QUBO)
Consider a graph $G_{p}=(V_{p},E_{p},w_{p})$ with node set $V_{p}=\{1,2,\ldots,n_{p}\}$, edge set $E_{p}$, and edge weights $w_{p}$ representing or inherent in a CO problem instance $p$. PI-GNN is a general UL framework for static CO problems based on QUBO (Lucas, 2014; Glover et al., 2018; Djidjev et al., 2018). A QUBO problem is defined by:
$\min_{x}~~\ell(x;Q_{p})=x^{T}Q_{p}x$ (1)
@@
MaxCut. Without loss of generality, we assume the edges have unit weights. The goal of MaxCut is to find a partition of the nodes into two disjoint subsets that maximizes the number of edges between them. We can model MaxCut solutions using $x\in\{0,1\}^{n}$ where $x_{i}=0$ if node $i$ is in one set, and $x_{i}=1$ if node $i$ is in the other set. The objective can be formulated as $\sum_{(i,j)\in E}(2x_{i}x_{j}-x_{i}-x_{j})$, which is an instance of QUBO (Glover et al., 2018).
MIS. MIS is the largest subset of non-adjacent nodes. Similar to MaxCut, we let $x_{i}=1$ if node $i$ is in the independent set and $x_{i}=0$ otherwise. The objective can be formulated as $\sum_{i\in V}x_{i}+M\sum_{(i,j)\in E}x_{i}x_{j}$, where $M\in\mathbb{R}_{>0}$ is a penalty term enforcing the independent set constraint (Djidjev et al., 2018).
TSP. We consider symmetric TSP instances modeled as undirected complete graphs. TSP aims to find the route that visits each node exactly once and returns to the origin node with the lowest total distance traveled. In contrast to MaxCut and MIS, we use a binary matrix $X\in\{0,1\}^{n\times n}$ to represent a route where $X_{ij}=1$ denotes visiting node $i$ at step $j$. Following Lucas (2014) and The MathWorks Inc. (2024), we adopt the following objective: $\sum_{(i,j)\in E}w_{ij}\sum_{v=1}^{n}X_{iv}X_{j(v+1)}+M\sum_{i=1}^{n}(1-\sum_{j=1}^{n}X_{ij})^{2}+M\sum_{j=1}^{n}(1-\sum_{i=1}^{n}X_{ij})^{2}$, where $w_{ij}$ is the distance between node $i$ and node $j$, $X_{j(n+1)}=X_{j1}$, accounting for the requirement of returning to the origin node, and $M\in\mathbb{R}_{>0}$ is a penalty term to ensure $X$ represents a valid route. During the QUBO objective computation, we simply flatten $X$ to an $n^{2}$-dimensional vector.
@@ 3.2 DCO on graphs
We focus on DCO problems on discrete-time dynamic graphs (DTDGs). A DTDG representing a DCO instance is a sequence $G_{p}^{1},G_{p}^{2},\ldots,G_{p}^{T}]$ of graph snapshots where each $G_{p}^{t}=(V_{p}^{t},E_{p}^{t},w_{p}^{t})$ has a node set $V_{p}^{t}=\{1,2,\ldots,n_{p}^{t}\}$, an edge set $E_{p}^{t}$, and edge weights $w_{p}^{t}$. For dynamic MaxCut and MIS, we consider the scenario where the node set $V_{p}^{t}$ and edge set $E_{p}^{t}$ change over time, and $w_{p}^{t}=\mathbf{1}$ for all $t\in\{1,2,\ldots,T\}$. For dynamic TSP, we consider the case where one of the nodes moves along a trajectory. Effectively, $V_{p}^{1}=V_{p}^{2}=\ldots=V_{p}^{T}$ and $E_{p}^{1}=E_{p}^{2}=\ldots=E_{p}^{T}$, but $w_{p}^{t}$ varies for each $t$.
Let $\Omega_{p}^{t}$ denote the set of discrete feasible solutions for snapshot $G_{p}^{t}$. For MaxCut and MIS, $\Omega_{p}^{t}$ is a subset of $\{0,1\}^{n_{p}^{t}}$. For TSP, $\Omega_{p}^{t}$ is the set of all possible routes that visit each node exactly once and return to the origin node. The objective is to find the optimal solution for each snapshot of a given DCO problem instance: $x^{t*}=\operatorname*{argmin}_{x\in\Omega_{p}^{t}}\ell(x;Q_{p}^{t})$, where $\ell(\cdot;Q_{p}^{t})$ denotes the cost function as defined earlier for each snapshot. For evaluation, we compute the mean approximation ratio (ApR) across all snapshots: $\text{Mean ApR}=\frac{1}{T}\sum_{t=1}^{T}\ell(x^{t};Q_{p}^{t})/\ell(x^{t*};Q_{p}^{t})$.
@@ 4.1 Warm-starting PI-GNN
Given a sequence of graph snapshots $G_{p}^{1},G_{p}^{2},\ldots,G_{p}^{T}]$, PI-GNN independently initiates an optimization process for each snapshot. One potential issue is that the time interval between two consecutive snapshots may be shorter than the convergence time of PI-GNN. In such cases, faster convergence is required. Considering the structural similarity across snapshots, a straightforward baseline method is to warm start the optimization using the parameters from the previous snapshot. Effectively, the “solution” to the previous snapshot serves as the initialization for the current one, allowing the model to “fine-tune” the solution instead of optimizing everything from scratch. We detail the procedure of warm-starting PI-GNN under different time budgets in Algorithm 1.
@@
2:Problem instance: $G_{p}^{1},G_{p}^{2},\ldots,G_{p}^{T}]$
@@
6:Randomly initialize $\theta^{1}$
7:Construct $Q_{p}^{1},Q_{p}^{2},\ldots,Q_{p}^{T}]$
8:for $G_{p}^{1},G_{p}^{2},\ldots,G_{p}^{T}]$
@@
10:  Predict $x^{1}$ via $\mathcal{A}(G_{p}^{1};\theta^{1})$
11:  Update $\theta^{1}$ using $\nabla\ell(x^{1};Q_{p}^{1})$
@@
14:  $\theta^{t}\leftarrow\theta^{t-1}$
@@
16:   Predict $x^{t}$ via $\mathcal{A}(G_{p}^{t};\theta^{t})$
17:   Update $\theta^{t}$ using $\nabla\ell(x^{t};Q_{p}^{t})$
@@
2:Problem instance: $G_{p}^{1},G_{p}^{2},\ldots,G_{p}^{T}]$
@@
6:Randomly initialize $\theta^{1}$
7:Construct $Q_{p}^{1},Q_{p}^{2},\ldots,Q_{p}^{T}]$
8:for $G_{p}^{1},G_{p}^{2},\ldots,G_{p}^{T}]$
@@
10:  Predict $x^{1}$ via $\mathcal{A}(G_{p}^{1};\theta^{1})$
11:  Update $\theta^{1}$ using $\nabla\ell(x^{1};Q_{p}^{1})$
@@
14:  $\theta^{t}\leftarrow\lambda_{\text{shrink}}\theta^{t-1}+\lambda_{\text{perturb}}\epsilon^{t}$
@@
16:   Predict $x^{t}$ via $\mathcal{A}(G_{p}^{t};\theta^{t})$
17:   Update $\theta^{t}$ using $\nabla\ell(x^{t};Q_{p}^{t})$
@@
Figure 1: Performance ($\uparrow$) of static PI-GNN and warm-started PI-GNN on dynamic MaxCut and MIS instances. Detailed setup is explained in Section 5.
These phenomena are closely tied to the role of initialization in deep learning (Sutskever et al., 2013; Glorot and Bengio, 2010; He et al., 2015) and can be attributed to the highly nonconvex nature of the optimization objective with respect to the model parameters, which gives rise to numerous local optima. When the first snapshot is optimized to convergence, the model undergoes extensive optimization, resulting in increased confidence in its predictions. Consequently, when the model is warm-started, the gradients become small due to this overconfidence, making it more difficult for the model to escape local optima. This hampers its ability to discover higher-quality solutions despite faster initial progress.